OpenStudy (anonymous):

The value 5π/24is a solution for the equation 4cos^2(4x)-3=0. A. True B. False

OpenStudy (anonymous):

\[4\cos ^{2}(4x)-3=0\]\[\cos ^{2}(4x)=3/4\]\[\cos (4x)=\sqrt{3}/2 or \cos (4x)=-\sqrt{3}/2\] either 4x=π/6 ----> x=π/24 , x=11π/24 or x = 5π/24 , x=7π/24 answer: A. True